OpenStudy (anonymous):

log_3[-x(x-2)]=log_3(x^2)

OpenStudy (anonymous):

Solve

OpenStudy (anonymous):

The answer is x=1 but I got x=0

OpenStudy (anonymous):

Can you write it using equation editor?

OpenStudy (anonymous):

\[\log_3[-x(x-2)]=\log_3(x^2)\]

myininaya (myininaya):

Did you set the insides = ?

myininaya (myininaya):

\[-x(x-2)=x^2 ? \]

OpenStudy (anonymous):